OpenStudy (anonymous):

find the interval of convergence of the following power series : [1*3*5*7*...*(2n+1)/4*9*14*...*(5n-1)] * (x-1)^n find the interval of convergence of the following power series : [1*3*5*7*...*(2n+1)/4*9*14*...*(5n-1)] * (x-1)^n @Mathematics

OpenStudy (anonymous):

This is a good question

OpenStudy (anonymous):

let me write it again first

OpenStudy (anonymous):

Infinite Series :O .. OMG !! I haven't studied this stuff yet.

OpenStudy (anonymous):

\[\sum_{n = 0}^{\infty}\frac{1*3*5*7*...*(2n+1)}{4*9*14*...*(5n-1)}(x-1)^n\]

OpenStudy (anonymous):

have you already found the radius of convergence?

OpenStudy (anonymous):

we can simplify the denominator a bit.

OpenStudy (anonymous):

Nope..to find the radius and interval of convergence i need the general term of the series

OpenStudy (anonymous):

the denominator is actually 2^n mutiplied by the sequence of odd numbers

OpenStudy (anonymous):

that way you can actually simplify the whole thing to \[\sum_{n=0}^{\infty}{\frac{1}{2^n}}(x-1)^n\]

OpenStudy (anonymous):

now we can use the root test!

OpenStudy (anonymous):

woah..awesome..thanks alot. but can you tell me how the denominator is 2^n * odd series? i dont see it :S

OpenStudy (anonymous):

yeah you're right.... I don't think it is 2^n * odd series

OpenStudy (anonymous):

hmm..cant figure it out

OpenStudy (anonymous):

lets just use the ratio test right away \[\lim_{n\rightarrow\infty}|\frac{a_{n+1}}{a_n}| = \lim_{n\rightarrow\infty} \frac{2n+3} {5n+4}*(x-1)\implies|x-1|<\frac{5}{2}\]

OpenStudy (anonymous):

but if we use the ratio test right away it only takes care of one term right? thought we need a general term..

OpenStudy (anonymous):

so our interval of convergence is \[-\frac{3}{2}<x<\frac{7}{2}\]

OpenStudy (anonymous):

nah we can divide all the terms out.

OpenStudy (anonymous):

I think we must test those limits now :-D

OpenStudy (anonymous):

lets try the left one, -3/2 this makes it \[\sum_{n = 0}^{\infty}\frac{1*3*5*7*...*(2n+1)}{4*9*14*...*(5n-1)}(-\frac{5}{2})^n\]

OpenStudy (anonymous):

hmm okay yeah now if its convergent at both the end points of the interval then its absolutely convergent, and if not it'll be conditionally covergent at one of them which is what i needed. thanks alot!

OpenStudy (anonymous):

let's look for a pattern.

OpenStudy (anonymous):

\[\sum_{n = 0}^{\infty}\frac{1*3*5*7*...*(2n+1)(5^n)}{4*9*14*...*(5n-1)(2^n)}(-1)^n\]=\[\sum_{n = 0}^{\infty}\frac{5*15*25*35*...*5(2n+1)}{8*18*28*...*2(5n-1)}(-1)^n\]

OpenStudy (anonymous):

be right back..